# Coherence-guided Preference Disentanglement for Cross-domain Recommendations

### v1

@@ Abstract.
Discovering user preferences across different domains is pivotal in cross-domain recommendation systems, particularly when platforms lack comprehensive user-item interactive data. The limited presence of shared users often hampers the effective modeling of common preferences. While leveraging shared items’ attributes, such as category and popularity, can enhance cross-domain recommendation performance, the scarcity of shared items between domains has limited research in this area. To address this, we propose a Coherence-guided Preference Disentanglement (CoPD) method aimed at improving cross-domain recommendation by i) explicitly extracting shared item attributes to guide the learning of shared user preferences and ii) disentangling these preferences to identify specific user interests transferred between domains. CoPD introduces coherence constraints on item embeddings of shared and specific domains, aiding in extracting shared attributes. Moreover, it utilizes these attributes to guide the disentanglement of user preferences into separate embeddings for interest and conformity through a popularity-weighted loss. Experiments conducted on real-world datasets demonstrate the superior performance of our proposed CoPD over existing competitive baselines, highlighting its effectiveness in enhancing cross-domain recommendation performance.
@@
copyright: acmlicensedjournalyear: 2024ccs: Information systems Recommender systems
@@ 1. Introduction
Cross-domain recommendation (CDR) (Tang et al., 2012; Li and Tuzhilin, 2019; Fernández-Tobías et al., 2012; Zhu et al., 2019; Cao et al., 2022; Liu et al., 2020) provides a strategy to alleviate the issues caused by sparse user-item interaction data, which generally utilize transfer learning approaches (Li and Tuzhilin, 2019; Fernández-Tobías et al., 2012) to transfer users’ preference information from a source domain to a target domain by employing their shared users. For example, when a user views a movie, the CDR methods can recommend a book with the same theme as the movie in the Book domain because the CDR methods can capture the shared users’ preferences. On the other hand, understanding a user’s preference for books facilitates us in recommending related movies. However, the scarcity of shared users across two domains significantly constrains the performance of cross-domain recommendations (Tang et al., 2012; Li and Tuzhilin, 2019). Tang et al. (Tang et al., 2012) emphasized the challenge of sparse overlapping users between domains, hindering the effective transfer of knowledge. Similarly, Li et al. (Li and She, 2017) highlighted domain shift and heterogeneity issues, where differences in item distributions and user behaviors across domains posed obstacles in aligning representations and transferring preferences. Fernandez et al. (Fernández-Tobías et al., 2012) noted that the cold-start problem, particularly affecting new user or item entries in cross-domain systems, exacerbating the challenge of accurate recommendations based on limited shared information. Recently, some dual-target CDR methods have successfully improved the results with sparse overlapping users (Zhu et al., 2019; Li and Tuzhilin, 2019; Zhu et al., 2020; Zhao et al., 2019). These methods integrate the information of different domains by combining the representations of overlapping users, establishing the connection between domains, and enhancing recommendation performance across multiple domains (Zang et al., 2022; Zhu et al., 2021a). Nevertheless, Zhu et al. (Zhu et al., 2019) identified a potential loss of domain-specific information during shared preference transfer, risking oversimplification and excluding crucial domain-specific characteristics that could enhance accuracy. To tackle this problem, some dual-target CDR methods aim to disentangle domain-shared and domain-specific information (Cao et al., 2022; Zhu et al., 2023), thereby improving model transfer effectiveness.
@@ 2.1. Cross-Domain Recommendations
The cross-domain recommendation aims to augment the recommendation effectiveness within the target domain by leveraging insights from other domains. Based on the inference goal of recommendation task, cross-domain recommendation methodologies can be roughly categorized into three categories (Liu et al., 2024): single-target domain (Rendle et al., 2012; He et al., 2017, 2020; Wang et al., 2019), dual-target domain (Liu et al., 2020; Cao et al., 2022; Zhu et al., 2023; Guo et al., 2022; Li and Tuzhilin, 2019; Zhu et al., 2020), and multi-target domain (Cui et al., 2020; Krishnan et al., 2020). Our CoPD belongs to dual-target cross-domain recommendations, focusing on transferring knowledge across two domains to improve recommendation performance in both domains. Dual-target CDR models emphasize identifying overlapping entities across disparate fields and achieving integration by sharing or enhancing the representations of such entities in distinct fields. Some approaches such as CDFM (Loni et al., 2014), employ shallow models as a backbone. CDFM aims to proficiently transfer knowledge by incorporating source domain interactions as supplementary features, based on the principles of Factorization Machines (Rendle, 2010). Furthermore, other approaches such as CoNet (Hu et al., 2018), utilize neural networks as a backbone. CoNet facilitates knowledge transfer across domains through the use of cross-connection networks between feed-forward multi-level perceptrons (MLPs). In a different view, DDTCDR (Li and Tuzhilin, 2019) pursues the learning of a latent orthogonal mapping function for shared users, enabling the transfer of user preferences across domains. Some methods leverage graph neural networks as a backbone, achieving state-of-the-art performance. For example, PPGN (Zhao et al., 2019) involves stacking several graph convolution networks on two domains and achieves knowledge transfer by sharing common hidden user features. Similarly, BiTGCF (Liu et al., 2020) introduced a graph collaborative filtering network, leveraging public users as conduits to facilitate bidirectional cross-domain knowledge transfer, thereby enhancing performance across both domains. DA-GCN (Guo et al., 2021) and TiDA-GCN (Guo et al., 2024) construct a Cross-Domain Sequence (CDS) graph to model complex associations among users and items across domains, where users and items from each domain are linked as nodes in a graph, with their associations represented as edges. Additionally, they are domain-aware in the sense that they adapt to the specific preferences and behaviors of users within each domain. However, previous methods neglect the disentanglement of domain-shared and domain-specific information, leading to the problem of negative transfer (Zhou et al., 2023; Zhu et al., 2021b). Wang et al. (Wang et al., 2022) utilizes the explicit reconstruction loss integrated within the Evidence Lower Bound (ELBO), complemented by additional regularizers such as the disentanglement loss, to learn the desirable disentangled representations. DisenCDR (Cao et al., 2022) propose a variational bipartite graph encoder (VBGE), and use different VBGEs encode domain-shared and domain-specific representations. Then, DisenCDR learns these representation with two disentanglement regularizers based on mutual information. DIDA-CDR (Zhu et al., 2023) captures the domain-shared user representations using an interpolative data augmentation approach and proposes a disentangling module to disentangle domain-specific and domain-independent information to capture comprehensive user preferences.
@@ 2.3. Differences
Our CoPD differs significantly from previous studies in several aspects: (1) Before disentangling user preference, we introduce a global-local graph encoder to disentangle domain-shared and domain-specific information. Different from DisenCDR (Cao et al., 2022) and DIDA-CDR (Zhu et al., 2023), CoPD learns domain-shared user representations through a global graph encoder, allowing information from both domains to influence the user representation. Additionally, we can learn domain-shared item representations using the global graph encoder and two coherence constraint regularizers, which have not been explored in previous dual-target CDR methods. Unlike PPGN (Zhao et al., 2019), which constructs a cross-domain user-item interaction graph to learn user and item representations, CoPD acquires domain-specific user representations through a local graph encoder. This approach allows CoPD to address the negative transfer problem and ensure independent learning of domain-specific representations. (2) Our preference disentanglement strategy differs from other disentangled representation learning methods. While previous methods aim to disentangle both user and item embeddings (Wang et al., 2020; Zheng et al., 2021), accomplishing this in dual-target CDR tasks presents challenges due to the data sparsity problem. Hence, CoPD exclusively focuses on disentangling user embeddings into interest and conformity embeddings and disentangling user and item embeddings into domain-shared and domain-specific embeddings. The learning of users’ interest and conformity embeddings is guided by disentangled domain-shared item embeddings. Specifically, to learn domain-shared user preferences, CoPD disentangles domain-shared user embeddings into domain-shared interest and conformity embeddings, which are guided by domain-shared item attribute embeddings. Furthermore, CoPD employs two preference encoders to disentangle preference embeddings. However, it is difficult to sample negative samples in CDR (Ma et al., 2023). To facilitate the effective training of these preference encoders, we employ a popularity-weighted Softmax loss, which enables the selection of challenging examples from multiple negative. Based on the causality analysis in DICE (Zheng et al., 2021), although user interests can indeed align with popular items, it is difficult to learn real interests from popular items. We assert an inverse correlation between interest embeddings and popularity weight to learn users’ real interests.
@@ 3.5. User Preference Disentanglement Guided by Aligned Item Attributes
After obtaining the disentangled preference embeddings, we hope $z_{u}^{int}$ and $z_{u}^{conf}$ reflect the preferences of user interest and conformity. We introduced the domain-shared item embeddings as mentioned in Section 3.4 (i.e., $z_{i}^{A,S}$ and $z_{i}^{B,S}$), which encompasses common attributes for domain A and domain B, to facilitate the representation learning for disentangled preference embeddings. Additionally, we design a popularity-weighted loss to control the impacts of different preferences. Specifically, Bayesian Personalized Ranking(BPR) loss (Rendle et al., 2012) and Softmax loss (Rendle, 2021) can be used to measure the quality of interaction predictions. We employ the Softmax loss here as it inherently facilitates the selection of challenging examples from multiple negative (Lian et al., 2020; Yi et al., 2019; Zhang et al., 2022, 2023), thus amplifying the process of learning particular user preference. There are two Softmax losses (i.e., conformity loss and interest loss) for learning the interest and conformity preferences. We aim to establish a positive correlation between conformity loss and popularity weight, as the model emphasizes conformity aspects for popular items. Conversely, we seek a negative correlation between interest loss and popularity weight, indicating that the model prioritizes capturing user interest over popularity for effective recommendations. We utilize item popularity as a measure of the popularity weight. Specifically, item popularity is determined by its popularity degree, which indicates that an item is more likely to be popular when it has been interacted with by many users. Taking domain A as an example, item popularity can be defined as $p_{i}^{A}$, it indicates the number of historical interactions involving the item, and we normalize $p_{i}^{A}$ range of 0 to 1 using min-max normalization. We minimize the popularity-weighted Softmax losses for representative learning:
@@ 4. Experiments
In this work, we claim that (1) the global graph encoder and coherence constraints can help extract domain-shared attributes for item embedding and (2) our CoPD can disentangle users’ interest and conformity embeddings guided by domain-shared item embedding. To verify the effectiveness of our CoPD, we conduct extensive experiments on four cross-domain recommendation tasks. Specifically, we provide comprehensive analyses from the subsequent perspectives:
@@
Table 2. Statistics of four dual-target CDR tasks.
@@ 4.1. Datasets
To evaluate the recommendation performance of our proposed CoPD and ensure a fair comparison with previous methods, we evaluate our model on four publicly available Amazon datasets [^1] http://jmcauley.ucsd.edu/data/amazon/ (He and McAuley, 2016; McAuley et al., 2015) and three Douban datasets(Zhu et al., 2020). The four Amazon datasets are Electronics (Elec), Cell Phones (Phone), Accessories, Sports and Outdoors (Sport) and Clothing, Shoes and Jewelry (Cloth). Following the method introduced in (Liu et al., 2020), we convert the explicit ratings into implicit feedback, representing ratings as 0 and 1 to indicate whether the user has rated the item. Next, we filter out users with fewer than 5 interactions and items with fewer than 10 interactions. The users in our dual-target cross-domain recommendation tasks are all overlapped, meaning the users at least click one item in each domain; thus, we extract overlapping users in both domains. Following the experimental setting in DisenCDR (Cao et al., 2022), we exclude cold-start items from the test set. Finally, we divide the aforementioned four datasets into four dual-target Cross-Domain Recommendation (CDR) tasks, comprising two tasks with high domain similarity (Elec & Phone, Sport & Cloth) and two tasks with lower domain similarity (Sport & Phone, Elec & Cloth). The three Douban datasets are Douban-Movie, Douban-Book and Douban-Music. Following the method introduced in GA-DTCDR (Zhu et al., 2020) and DIDA-CDR (Zhu et al., 2023). We convert the explicit ratings into implicit feedback. Next, we filter out users and items with fewer than 5 interactions. We also filter out of the cold-start item entry in the test set. Finally, we construct the three Douban datasets for two dual-target CDR tasks. The details of these four six dual-target CDR tasks are presented in Table 2.
@@ 4.2. Evaluation Metrics
In item recommendation, the Leave-One-Out (LOO) evaluation (He et al., 2016) is commonly employed to demonstrate performance (Cao et al., 2022; Liu et al., 2020; Hu et al., 2018; He et al., 2017). To compare the performance of CoPD and other baseline models, we adopt the LOO evaluation method. To construct a test set of each user, one item is preserved as ground truth, and we randomly sample 999 items that the user has not interacted with as negative items. We utilize two widely adopted evaluation metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG) (Järvelin and Kekäläinen, 2002). The predicted rank list is cut off at $K$ = 10. The HR measures whether the test items are on the top-10 list, and the NDCG estimates ranking quality by assigning higher scores to hits at the top ranks. In both metrics, a higher value indicates superior recommendation performance.

### v2

@@ Abstract.
Discovering user preferences across different domains is pivotal in cross-domain recommendation systems, particularly when platforms lack comprehensive user-item interactive data. The limited presence of shared users often hampers the effective modeling of common preferences. While leveraging shared items’ attributes, such as category and popularity, can enhance cross-domain recommendation performance, the scarcity of shared items between domains has limited research in this area. To address this, we propose a Coherence-guided Preference Disentanglement (CoPD) method aimed at improving cross-domain recommendation by i) explicitly extracting shared item attributes to guide the learning of shared user preferences and ii) disentangling these preferences to identify specific user interests transferred between domains. CoPD introduces coherence constraints on item embeddings of shared and specific domains, aiding in extracting shared attributes. Moreover, it utilizes these attributes to guide the disentanglement of user preferences into separate embeddings for interest and conformity through a popularity-weighted loss. Experiments conducted on real-world datasets demonstrate the superior performance of our proposed CoPD over existing competitive baselines, highlighting its effectiveness in enhancing cross-domain recommendation performance. The code is available at https://github.com/XiangZongyi/CoPD.
@@
copyright: acmlicensedjournalyear: 2025ccs: Information systems Recommender systems
@@ 1. Introduction
Cross-domain recommendation (CDR) (Tang et al., 2012; Li and Tuzhilin, 2019; Fernández-Tobías et al., 2012; Zhu et al., 2019; Cao et al., 2022; Liu et al., 2020; Xu et al., 2022) provides a strategy to alleviate the issues caused by sparse user-item interaction data, which generally utilize transfer learning approaches (Li and Tuzhilin, 2019; Fernández-Tobías et al., 2012; Hu et al., 2022) to transfer users’ preference information from a source domain to a target domain by employing their shared users. For example, when a user views a movie, the CDR methods can recommend a book with the same theme as the movie in the Book domain because the CDR methods can capture the shared users’ preferences. On the other hand, understanding a user’s preference for books facilitates us in recommending related movies. However, the scarcity of shared users across two domains significantly constrains the performance of cross-domain recommendations (Tang et al., 2012; Li and Tuzhilin, 2019). Tang et al. (Tang et al., 2012) emphasized the challenge of sparse overlapping users between domains, hindering the effective transfer of knowledge. Similarly, Li et al. (Li and She, 2017) highlighted domain shift and heterogeneity issues, where differences in item distributions and user behaviors across domains posed obstacles in aligning representations and transferring preferences. Fernandez et al. (Fernández-Tobías et al., 2012) noted that the cold-start problem, particularly affecting new user or item entries in cross-domain systems, exacerbating the challenge of accurate recommendations based on limited shared information. Recently, some dual-target CDR methods have successfully improved the results with sparse overlapping users (Zhu et al., 2019; Li and Tuzhilin, 2019; Zhu et al., 2020; Zhao et al., 2019). These methods integrate the information of different domains by combining the representations of overlapping users, establishing the connection between domains, and enhancing recommendation performance across multiple domains (Zang et al., 2022; Zhu et al., 2021b). Nevertheless, Zhu et al. (Zhu et al., 2019) identified a potential loss of domain-specific information during shared preference transfer, risking oversimplification and excluding crucial domain-specific characteristics that could enhance accuracy. To tackle this problem, some dual-target CDR methods aim to disentangle domain-shared and domain-specific information (Cao et al., 2022; Zhu et al., 2023), thereby improving model transfer effectiveness.
@@ 2.1. Cross-Domain Recommendations
The cross-domain recommendation aims to augment the recommendation effectiveness within the target domain by leveraging insights from other domains. Based on the inference goal of recommendation task, cross-domain recommendation methodologies can be roughly categorized into three categories (Liu et al., 2024): single-target domain (Man et al., 2017; Zhu et al., 2021a), dual-target domain (Liu et al., 2020; Cao et al., 2022; Zhu et al., 2023; Guo et al., 2022; Li and Tuzhilin, 2019; Zhu et al., 2020), and multi-target domain (Cui et al., 2020; Krishnan et al., 2020). Our CoPD belongs to dual-target cross-domain recommendations, focusing on transferring knowledge across two domains to improve recommendation performance in both domains. Dual-target CDR models emphasize identifying overlapping entities across disparate fields and achieving integration by sharing or enhancing the representations of such entities in distinct fields. Some approaches such as CDFM (Loni et al., 2014), employ shallow models as a backbone. CDFM aims to proficiently transfer knowledge by incorporating source domain interactions as supplementary features, based on the principles of Factorization Machines (Rendle, 2010). Furthermore, other approaches such as CoNet (Hu et al., 2018), utilize neural networks as a backbone. CoNet facilitates knowledge transfer across domains through the use of cross-connection networks between feed-forward multi-level perceptrons (MLPs). In a different view, DDTCDR (Li and Tuzhilin, 2019) pursues the learning of a latent orthogonal mapping function for shared users, enabling the transfer of user preferences across domains. Some methods leverage graph neural networks as a backbone, achieving state-of-the-art performance. For example, PPGN (Zhao et al., 2019) involves stacking several graph convolution networks on two domains and achieves knowledge transfer by sharing common hidden user features. Similarly, BiTGCF (Liu et al., 2020) introduced a graph collaborative filtering network, leveraging public users as conduits to facilitate bidirectional cross-domain knowledge transfer, thereby enhancing performance across both domains. DA-GCN (Guo et al., 2021) and TiDA-GCN (Guo et al., 2024) construct a Cross-Domain Sequence (CDS) graph to model complex associations among users and items across domains, where users and items from each domain are linked as nodes in a graph, with their associations represented as edges. Additionally, they are domain-aware in the sense that they adapt to the specific preferences and behaviors of users within each domain. However, previous methods neglect the disentanglement of domain-shared and domain-specific information, leading to the problem of negative transfer (Zhou et al., 2023; Zhu et al., 2021c). Wang et al. (Wang et al., 2022) utilizes the explicit reconstruction loss integrated within the Evidence Lower Bound (ELBO), complemented by additional regularizers such as the disentanglement loss, to learn the desirable disentangled representations. DisenCDR (Cao et al., 2022) propose a variational bipartite graph encoder (VBGE), and use different VBGEs encode domain-shared and domain-specific representations. Then, DisenCDR learns these representation with two disentanglement regularizers based on mutual information. DIDA-CDR (Zhu et al., 2023) captures the domain-shared user representations using an interpolative data augmentation approach and proposes a disentangling module to disentangle domain-specific and domain-independent information to capture comprehensive user preferences.
@@ 2.3. Differences
Our CoPD differs significantly from previous studies in several aspects: (1) Before disentangling user preference, we introduce a global-local graph encoder to disentangle domain-shared and domain-specific information. Different from DisenCDR (Cao et al., 2022) and DIDA-CDR (Zhu et al., 2023), CoPD learns domain-shared user representations through a global graph encoder, allowing information from both domains to influence the user representation. Additionally, we can learn domain-shared item representations using the global graph encoder and two coherence constraint regularizers, which have not been explored in previous dual-target CDR methods. Unlike PPGN (Zhao et al., 2019), which constructs a cross-domain user-item interaction graph to learn user and item representations, CoPD acquires domain-specific user representations through a local graph encoder. This approach allows CoPD to address the negative transfer problem and ensure independent learning of domain-specific representations. (2) Our preference disentanglement strategy differs from other disentangled representation learning methods. While previous methods aim to disentangle both user and item embeddings (Wang et al., 2020; Zheng et al., 2021), accomplishing this in dual-target CDR tasks presents challenges due to the data sparsity problem. Hence, CoPD focuses on disentangling user embeddings into interest and conformity embeddings and disentangling user and item embeddings into domain-shared and domain-specific embeddings. The learning of users’ interest and conformity embeddings is guided by disentangled domain-shared item embeddings. Specifically, to learn domain-shared user preferences, CoPD disentangles domain-shared user embeddings into domain-shared interest and conformity embeddings, which are guided by domain-shared item attribute embeddings. Furthermore, CoPD employs two preference encoders to disentangle preference embeddings. However, it is difficult to sample negative samples in CDR (Ma et al., 2023). To facilitate the effective training of these preference encoders, we employ a popularity-weighted Softmax loss, which enables the selection of challenging examples from multiple negative. Based on the causality analysis in DICE (Zheng et al., 2021), although user interests can indeed align with popular items, it is difficult to learn real interests from popular items. We assert an inverse correlation between interest embeddings and popularity weight to learn users’ real interests.
@@ 3.5. User Preference Disentanglement Guided by Aligned Item Attributes
After obtaining the disentangled preference embeddings, we hope $z_{u}^{int}$ and $z_{u}^{conf}$ reflect the preferences of user interest and conformity. We introduced the domain-shared item embeddings as mentioned in Section 3.4 (i.e., $z_{i}^{A,S}$ and $z_{i}^{B,S}$), which encompasses common attributes for domain A and domain B, to facilitate the representation learning for disentangled preference embeddings. Bayesian Personalized Ranking(BPR) loss (Rendle et al., 2012) and Softmax loss (Rendle, 2021) can be used to measure the quality of interaction predictions. We employ the Softmax loss here as it inherently facilitates the selection of challenging examples from multiple negative (Lian et al., 2020; Yi et al., 2019; Zhang et al., 2022, 2023), thus amplifying the process of learning particular user preference. There are two Softmax losses (i.e., conformity loss $\mathcal{L}_{conf}$ and interest loss $\mathcal{L}_{int}$) for learning the interest and conformity preferences. However, if the losses are simply summed without distinction for different items, the leaned representation $z_{u}^{int}$ and $z_{u}^{conf}$ cannot accurately reflect the preferences for user interest and conformity over items. Therefore, we integrate item popularity signal as a weight into the Softmax losses. Item popularity $p_{i}$ is defined as the number of historical interactions that an item $i$ has been involved in. A more popular item $i$ has a higher item popularity $p_{i}$. We then normalize $p_{i}$ to a range of 0 to 1 using min-max normalization. We aim to establish a positive correlation between the conformity loss $\mathcal{L}_{conf}$ and the popularity weight $p_{i}$. For highly popular items, the weight $p_{i}$ in the conformity $\mathcal{L}_{conf}$ would be close to 1, ensuring that popular items are primarily attributed to user’s conformity preference $z_{u}^{conf}$. Conversely, we aim to establish a negative correlation between the interest loss $\mathcal{L}_{int}$ and the popularity weight $p_{i}$. We use $1-p_{i}$ as the weight of $\mathcal{L}_{int}$. For less popular items, the weight $1-p_{i}$ would be close to 1, allowing the model to capture the users’ interest preference $z_{u}^{int}$ for certain item attributes. For more popular items, the weight would be close to 0 , thereby reducing the influence of item popularity and enabling the model to learn the user’s genuine interests. Taking domain A as an example, popularity weight in Softmax loss can be defined as $p_{i}^{A}$. We minimize the popularity-weighted Softmax losses for representative learning:
@@ 4. Experiments
In this work, we claim that (1) the global graph encoder and coherence constraints can help extract domain-shared attributes for item embedding and (2) our CoPD can disentangle users’ interest and conformity embeddings guided by domain-shared item embedding. To verify the effectiveness of our CoPD, we conduct extensive experiments on six cross-domain recommendation tasks. Specifically, we provide comprehensive analyses from the subsequent perspectives:
@@
Table 2. Statistics of six dual-target CDR tasks.
@@ 4.1. Datasets
To evaluate the recommendation performance of our proposed CoPD and ensure a fair comparison with previous methods, we evaluate our model on four publicly available Amazon datasets [^1] http://jmcauley.ucsd.edu/data/amazon/ (He and McAuley, 2016; McAuley et al., 2015) and three Douban datasets(Zhu et al., 2020). The four Amazon datasets are Electronics (Elec), Cell Phones (Phone), Accessories, Sports and Outdoors (Sport) and Clothing, Shoes and Jewelry (Cloth). Following the method introduced in (Liu et al., 2020), we convert the explicit ratings into implicit feedback, representing ratings as 0 and 1 to indicate whether the user has rated the item. Next, we filter out users with fewer than 5 interactions and items with fewer than 10 interactions. The users in our dual-target cross-domain recommendation tasks are all overlapped, meaning the users at least click one item in each domain; thus, we extract overlapping users in both domains. Following the experimental setting in DisenCDR (Cao et al., 2022), we exclude cold-start items from the test set. Finally, we divide the aforementioned four Amazon datasets into four dual-target Cross-Domain Recommendation (CDR) tasks, comprising two tasks with high domain similarity (Elec & Phone, Sport & Cloth) and two tasks with lower domain similarity (Sport & Phone, Elec & Cloth). The three Douban datasets are Douban-Movie, Douban-Book and Douban-Music. Following the method introduced in GA-DTCDR (Zhu et al., 2020) and DIDA-CDR (Zhu et al., 2023). We convert the explicit ratings into implicit feedback. Next, we filter out users and items with fewer than 5 interactions. We also filter out of the cold-start item entry in the test set. Finally, we construct the three Douban datasets for two dual-target CDR tasks. The details of these six dual-target CDR tasks are presented in Table 2.
@@ 4.2. Evaluation Metrics
In item recommendation, the Leave-One-Out (LOO) evaluation (He et al., 2016) is commonly employed to demonstrate performance (Cao et al., 2022; Liu et al., 2020; Hu et al., 2018; He et al., 2017; Wu et al., 2023). To compare the performance of CoPD and other baseline models, we adopt the LOO evaluation method. To construct a test set of each user, one item is preserved as ground truth, and we randomly sample 999 items that the user has not interacted with as negative items. We utilize two widely adopted evaluation metrics: Hit Ratio (HR) and Normalized Discounted Cumulative Gain (NDCG) (Järvelin and Kekäläinen, 2002). The predicted rank list is cut off at $K$ = 10. The HR measures whether the test items are on the top-10 list, and the NDCG estimates ranking quality by assigning higher scores to hits at the top ranks. In both metrics, a higher value indicates superior recommendation performance.